OpenStudy (anonymous):

How many solutions does the equation have 2(2X+5)=4(X+3) A.one solution B infinite solutions C no solution

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

2(2x +5) =4(x+3) doing the same thing earlier you'll have 4x + 10 = 4x + 12 4x - 4x = 12 - 10 0 is not equal to 2 so what do you think is the answer?

OpenStudy (anonymous):

b @shinalcantara

OpenStudy (shinalcantara):

remember that when they are not equal like this, there's no solution and when they are equal then it has infinite solution. if it had been that the equation resulted to 4x+10=4x+10 then you'll have the solution as infinite but the case is different. the answer should be C

OpenStudy (anonymous):

thank you so much I am not good at this @shinalcantara

OpenStudy (shinalcantara):

yw ;)